# Gravitation and Spacetime: Emergent from Spinor Interactions - How?

### v2

@@ 6.4 Symplectic Structure and Weyl Algebra
For a graph $\Gamma$ with edges intersecting the boundary $S$ of a spatial section $\Sigma_{0}$ in a causal double cone $\mathcal{D}$, we can define a Weyl algebra.
A (spin network) state $\omega$ over the algebra ${\cal A}(\Sigma)$ may be defined by representations on a closed, oriented differentiable graph $\Gamma$ embedded in $\Sigma_{0}$, with differentiable edges $e\in E$ intersected transversally by a differentiable $2$-dimensional oriented surface $S$ at a countable of intersection vertices $v\in V$. Let $C_{\Gamma}\in\mbox{\rm Cyl}$ be a $\mathcal{C}$$^{\infty}$ Cylinder function with respect to an $\mbox{\rm SO}(d)$ holonomy group on $\Gamma$, i.e. on each closed finite subgraph $\Gamma^{\prime}\subset\Gamma$ it is $C_{\Gamma^{\prime}}:=c(g_{1},\ldots,g_{N})$ where $g_{k}\in\mbox{\rm SO}(d)$, and $c$ is a differentiable function. With test function $f$ a derivation $X_{S,f}$ on Cyl is defined by
@@
Here the classical (extended) phase space is the cotangent bundle $T^{*}{\mathcal{C}}$ over a space $\mathcal{A}$ of (suitably regular) finitely localized connections. Let $\delta=({\delta_{A}},{\delta_{E}})\in T^{*}{\mathcal{C}}$. With suitable boundary conditions, a (weakly non-degenerate) symplectic form $\Omega$ over $T^{*}{\mathcal{C}}$ acts via
@@
with normalizing scale factor $\ell$. After lifting from $\mathcal{C}$ to $T^{*}{\mathcal{C}}$, the cylinder functions $q\in\mbox{\rm Cyl}$ serve as (gauge invariant) classical elementary configuration functions on $T^{*}{\mathcal{C}}$. The derivations $p\in\mbox{\rm Der}(\mbox{\rm Cyl})$ serve as classical elementary momentum functions on $T^{*}{\mathcal{C}}$. They are obtained as the Poisson-Lie action of $2$-dimensionally smeared duals of densitized triads $E$. $\mbox{\rm Cyl}\times\mbox{\rm Der}(\mbox{\rm Cyl})$ has a Poisson-Lie structure
@@
$\Omega_{0}\left({(\delta_{q},\delta_{p})},{(\delta_{q}^{\prime},\delta_{p}^{\prime})}\right)$ $:=$ $\int_{{\mathcal{C}_{\gamma\cup\gamma^{\prime}}}/{\mathcal{G}_{\gamma\cup\gamma^{\prime}}}}d\mu_{\gamma\cup\gamma^{\prime}}\left[pq^{\prime}-p^{\prime}q\right],$ (62)
@@
With symplectic form $\Omega$ to any function $\Omega(f,\cdot)$ a selfadjoint operator $\hat{\Omega}(f,\cdot)$ and a corresponding unitary Weyl element $W(f):=e^{i\hat{\Omega}(f,\cdot)}$ can be constructed, on the extension of an Hilbert space. With multiplication $W(f_{1})W(f_{2}):=e^{i\Omega(f_{1},f_{2})}W(f_{1}+f_{2})$, and conjugation $W(f)\mapsto W(-f)$, the Weyl elements generate a *-algebra. A norm on is defined by $\|f\|:=\frac{1}{4}\sup_{g\neq 0}\frac{\Omega(f,g)}{<g,g>}$. The $C^{*}$-closure under the sup-norm then generates a $C^{*}$-algebra.
In a gauge and $3$-diffeomorphism invariant representation of ${\cal A}_{\Gamma}$, typical observables in are represented by configuration multiplication operators $C_{\Gamma}\in\mbox{\rm Cyl}$ on Hilbert space $\mathcal{H}_{\Gamma}$, and by gauge-invariant and 3-diffeomorphism invariant combinations of derivative operators $X_{S,f}\in\mbox{\rm Der}(\mbox{\rm Cyl})$. Certain quadratic combinations e.g. yield the area operator.
Using holonomy flux variables, for each edge $e$ of $\Gamma$ intersecting the surface $S$, define
$U_{e}(\lambda)$ $=e^{i\lambda\hat{h}_{e}}$ (63)
$V_{e}(\mu)$ $=e^{i\mu\hat{E}_{e}}.$ (64)
These satisfy the Weyl relations
$U_{e}(\lambda)V_{e}(\mu)=e^{i\lambda\mu}V_{e}(\mu)U_{e}(\lambda).$ (65)
Operator products from all intersections generate the full Weyl algebra $\mathcal{W}(\Gamma,S)$ for the graph $\Gamma$ on the boundary surface $S$.
@@ 7 Conclusions
We have reviewed approaches to spacetime geometry in terms of spinor variables.
In any case, spinors are more fundamental than metric geometry:
@@
The sketchy developments in the current paper can be understood as proposal for further research along serveral open questions and into different possible directions.
@@
Massive particle modes, on the other hand, provide natural Compton clocks defining the the time and length scale of geometry, breaking the conformal invariance. The Higgs mechanism on the particle interaction network should be understood in more detail, and in its projection to spin networks. Similarly the coupling of matter to classical geometry should be recovered from the particle spinor interaction network.
@@
We argue, that the former will be a much more natural and efficient choice: The spin networks should arise as entanglement networks of elementary particle interactions. The latter being ultimately quantum in nature, the induced spin networks will inherit the quantum nature from the particle content of the spacetime vaccuum. That particular also explains the apparent similarity of classical Newtonian gravitational action at distance to quantum entanglement.
@@
Spacetime itself is a very successful concept, but emergent only at scales above the Planck length. From an atomistic perspective, its geometry must be derived from fundamental microscopic constituents. We have proposed that, a causal interaction network of all the elementary spinor particles ever created or anihilated in a defined causal domain may, together with their interactions, naturally generate the spin network which then provides their underlying UV-regularized space(time) geometry.

### v3

@@ 6.4 Symplectic Structure and Weyl Algebra
For a graph $\Gamma$ with edges intersecting the boundary $S$ of a spatial section $\Sigma_{0}$ in a causal double cone $\mathcal{D}$, we can define a Weyl algebra of observables localized thereon.
A (spin network) state $\omega$ over the algebra ${\cal A}(\Sigma_{0})$ may be defined by representations on a closed, oriented differentiable finite graph $\gamma$ embedded in $\Sigma_{0}$, with differentiable edges $e\in E$ intersected transversally by a differentiable $2$-dimensional oriented surface $S$ at a countable of intersection vertices $v\in V$. Let $C_{\Gamma}\in\mbox{\rm Cyl}$ be a $\mathcal{C}$$^{\infty}$ Cylinder function with respect to the holonomy group $G$ on $\Gamma$, i.e. $C_{\Gamma}:=c(g_{1},\ldots,g_{N})$ where $g_{k}\in G$, and $c$ is a differentiable function. With test function $f$ a derivation $X_{S,f}$ on Cyl is defined by
@@
Here the classical (extended) phase space is the cotangent bundle $P:=T^{*}{\mathcal{C}}$ over a space $\mathcal{A}$ of (suitably regular) finitely localized connections. Let $\delta=({\delta_{A}},{\delta_{E}})\in T^{*}{\mathcal{C}}$. With suitable boundary conditions, a (weakly non-degenerate) symplectic form $\Omega$ over $T_{e}P$ acts via
@@
with normalizing scale factor $\ell$. After lifting from $\mathcal{C}$ to phase space $P=T^{*}{\mathcal{C}}$, the cylinder functions $q\in\mbox{\rm Cyl}$ serve as (gauge invariant) configuration variables and the derivations $p\in\mbox{\rm Der}(\mbox{\rm Cyl})$ serve as conjugate momentum variables. They are obtained by the Poisson-Lie action of $2$-dimensionally smeared duals of densitized triads $E$. $\mbox{\rm Cyl}\times\mbox{\rm Der}(\mbox{\rm Cyl})$ has a Poisson-Lie structure
@@
$\Omega\left({(\delta_{q},\delta_{p})},{(\delta_{q}^{\prime},\delta_{p}^{\prime})}\right)$ $:=$ $\int_{{\mathcal{C}_{\gamma\cup\gamma^{\prime}}}/{\mathcal{G}_{\gamma\cup\gamma^{\prime}}}}d\mu_{\gamma\cup\gamma^{\prime}}\left[pq^{\prime}-p^{\prime}q\right],$ (62)
@@
With a symplectic form $\Omega$, quantization of a function $\Omega(f,\cdot)$ yields a selfadjoint operator $\hat{\Omega}(f,\cdot)$ and a corresponding unitary Weyl element $W(f):=e^{i\hat{\Omega}(f,\cdot)}$ can be constructed on a Hilbert space extension. With multiplication $W(f_{1})W(f_{2}):=e^{i\Omega(f_{1},f_{2})}W(f_{1}+f_{2})$, and conjugation $W(f)\mapsto W(-f)$, the Weyl elements generate a *-algebra. A norm on is defined by $\|f\|:=\frac{1}{4}\sup_{g\neq 0}\frac{\Omega(f,g)}{<g,g>}$. The $C^{*}$-closure under the operator sup-norm then generates a $C^{*}$-algebra from the Weyl elements.
In a gauge and $3$-diffeomorphism invariant representation of the algebra ${\cal A}_{\gamma}$ of observables localized at $\gamma$, the observables in are represented by configuration multiplication operators $C_{\gamma}\in\mbox{\rm Cyl}$ on Hilbert space $\mathcal{H}_{\gamma}$, and by gauge-invariant and 3-diffeomorphism invariant combinations with derivative operators $X_{S,f}\in\mbox{\rm Der}(\mbox{\rm Cyl})$. E.g., certain quadratic combinations yield the area operator with spectrum as in (32).
Alternatively, using holonomy flux variables $(h_{e}[A],E_{S}^{a}[E])$, for each edge $e$ of $\gamma$ intersecting the surface $S$ at puncture $p=e\cap S$, quantizations yield unitary operators
$U_{e}(\lambda)$ $=\hat{h}_{e}[\lambda A]$ (63)
$V_{e}(\mu)$ $=e^{i\mu\hat{E}_{e}}\quad,$ (64)
which do not commute at $p$ (cf. equations (28), (30), (31) above). Operator products from all intersection points $p$ then generate an algebra $\mathcal{W}(\gamma,S)$ for the graph $\gamma\subset\Sigma_{0}$ on the boundary surface $S=\partial\Sigma_{0}$.
@@ 7 Conclusions
We have reviewed approaches to spacetime geometry in terms of spinor variables. In any case, spinors are more fundamental than metric geometry:
@@
The sketchy developments in the current paper can be understood as proposal for further research along several open questions and into different possible directions.
@@
Massive particle modes, on the other hand, provide natural Compton clocks defining the time and length scale of geometry, breaking the conformal invariance. The Higgs mechanism on the particle interaction network should be understood in more detail, and in its projection to spin networks. Similarly the coupling of matter to classical geometry should be recovered from the particle spinor interaction network.
@@
We argue, that the former will be a much more natural and efficient choice: The spin networks should arise as entanglement networks of elementary particle interactions. The latter being ultimately quantum in nature, the induced spin networks will inherit the quantum nature from the particle content of the spacetime vacuum. That particular also explains the apparent similarity of classical Newtonian gravitational action at distance to quantum entanglement.
@@
Spacetime itself is a very successful concept, but emergent only at scales above the Planck length. From an atomistic perspective, its geometry must be derived from fundamental microscopic constituents. We have proposed that, a causal interaction network of all the elementary spinor particles ever created or annihilated in a defined causal domain may, together with their interactions, naturally generate the spin network which then provides their underlying UV-regularized space(time) geometry.